-5x+24=6-X
(the -5x+24 is in the radical)

Answer by prabhjyot(165)

sqrt(-5x+24)=6-X
Squaring both sides
-5x +24 = (6-x)^2
-5x =24 =x^2 +36 -12x
x^2- 7x +12=0
Now solving for x
step 1:Write down all the different ways to factor the constant coefficient (12) into two negative numbers:
12 = ( -1)(-12 )
12 = ( -2)( -6 )
12 = ( -3)( -4 )
(Order doesn’t matter. So (-3)(-4) is the same as (-4)(- 3).)
Step 2: Now, for each possible factorization, add together the two factors. You’re looking for two that add up to the middle coefficient, which is –7.
-1 – 12 = –13
-2 – 6 = –8
-3 – 4 = – 7 < --------------------------- found it!
The factors which worked out were –3 and –4, so here’s how
you factor the polynomial:
(x – 3)(x – 4)
So the solution to the equation is x=3,4